# SoRA: Singular Value Decomposed Low-Rank Adaptation for Domain Generalizable Representation Learning

### v1

@@ 5.4 Subject Personalization
DreamBooth (DB) [70] introduces subject personalization, which fine-tunes a pre-trained Text-to-Image (T2I) model on a few reference images to enable it to generate new visual concepts. DB updates all parameters within specific blocks, potentially impairing the pre-trained capability to generate images across various domains. Therefore, we explore whether SoRA can be seamlessly integrated into the DB framework to preserve pre-trained knowledge. For this task, we utilize Stable Diffusion XL (SDXL) [64] for the T2I diffusion model. Fig. 5 presents a visual comparison between DB and SoRA, using identical hyperparameter settings and sample seeds. As seen, SoRA, similar to the original SDXL model, can generate high-quality images of diverse visual domains (e.g., style, texture, pose) and exhibits superior identity preservation compared to DB with prior preservation (p.p.) loss [70]. More results with varied prompts and subjects are provided in the Supplemental.
@@ 6 Conclusion
This paper presents a multifaceted exploration of the distribution of generalizable components in VFMs—examining them at the weight level, block level, and within training dynamics—and introduces the SoRA framework as an effective means to preserve generalizable components. Contrary to recent approaches that passively refine frozen VFM features, SoRA is integrated into all linear layers to effectively facilitate task adaptation, while minimizing interference with generalizable components by initializing the low-rank adapter with minor singular components. Through extensive experiments, we observe that SoRA significantly outperforms the state-of-the-art methods in DGSS, DGOD, and even subject-driven generation tasks.
@@
For a comprehensive understanding of our proposed SoRA framework, we have provided this supplementary material. The following table of contents gives a concise overview and directs readers to specific sections of interest.
@@ A.1 DGSS Settings
The experimental settings for all studies conducted in the main paper are outlined in Tab. 11. Unless otherwise specified, Mask2Former [13] is utilized as the default decode head, and following Rein [80], we adopt only the basic data augmentation used in Mask2Former. Additionally, EMA is selectively employed to ensure stable training.
@@ A.2 DGOD Settings
The DGOD settings are detailed in the rightmost two columns of Tab. 11. When applying SoRA to convolution-based backbones such as ResNet [30], we linearize both the patch-level convolution and its weights. Specifically, a single convolution operation can be represented as a linear layer, , where , , and . We then apply SoRA as described in Eq. 1. For ResNet backbones, which possess a much narrower pre-trained knowledge compared to VFMs, we use extensive image corruption techniques to simulate domain shifts, following DivAlign [17]. In contrast, when using DINOv2 [58] as the backbone, we simply utilize basic data augmentation used in Co-DETR [93].
@@ A.3 Subject Personalization Settings
We conduct experiments on the DreamBooth dataset [70], which consists of 30 subjects with 4–6 images per subject. In all experiments the SoRA weights are trained using Adam optimizer for 500 iterations with a learning rate of . We set the adapter rank to and only use a center crop for data augmentation. Inspired by recent findings [25] that the first 10 attention layers of up_blocks.0 in SDXL [64] are pivotal for preserving image content, we fine-tune only these layers. Furthermore, to fully leverage pre-trained image-text joint representations, we freeze the cross-attention modules and apply SoRA solely to the self-attention modules.
Appendix B Detailed Ablations
B.1 Component Analysis
In this subsection, we conduct detailed ablation studies under multiple settings: GTAV Mapillary DGSS and Daytime-Sunny {Dusk-Rainy, Daytime-Foggy} DGOD scenarios. In Tables 13 and 13, we systematically evaluate the effectiveness of each component within the SoRA framework based on class-wise IoU/AP (%). All proposed components enhance overall generalization performance without adding any additional training or inference costs.
As illustrated in Tab. 13, freezing early blocks not only substantially reduces the number of trainable parameters but also significantly improves performance for classes that are infrequently observed in the source dataset (e.g., bicycle, motorcycle, train). Additionally, tuning minor singular components maximizes the retention of VFM’s world knowledge during task adaptation, leading to superior generalization performance over tuning principal components for most classes. Lastly, annealing weight decay proves especially beneficial for classes requiring fine-detail discrimination (e.g., road vs. sidewalk, traffic light vs. traffic sign, car vs. truck vs. bus vs. train, motorcycle vs. bicycle). Likewise, all components clearly improve recognition for the majority of classes under adverse weather conditions (see Tab. 13).
@@ B.2 Freezing Scheme
While freezing the initial blocks of VFM is effective in preserving its generalization ability during task adaptation, freezing too many blocks can lead to a reduction in discriminability. To better understand this trade-off, we explore the effects of varying the number of frozen blocks. Tab. 14 shows that freezing up to the first 8 blocks progressively enhances performance, but freezing beyond this point results in a decline. Considering that feature maps from multiple blocks (e.g., the 8th, 12th, 16th, and 24th blocks in large-sized backbones) serve as inputs to the segmentation/detection head, using more than one frozen VFM features as head input significantly undermines task adaptability (i.e. discriminability). Furthermore, since the first input feature map of the decode head is directly incorporated into the final mask prediction in Mask2Former [13], freezing the blocks that generate this feature map allows the full utilization of the generalization capacity of the early blocks in VFMs (see Fig. 3).
@@ C.1 Results on Various Backbones
Tab. 16 showcases the versatility of SoRA across a wide range of backbones, ranging from isotropic Vision Transformers (ViTs) to ConvNets and hierarchical ViT, as well as models trained under various approaches, such as ImageNet [18] supervision and MAE [31, 81] pre-training. SoRA consistently outperforms FFT across diverse backbone architectures. Notably, the improvements brought by SoRA become increasingly pronounced with larger model sizes and more extensive, high-quality data during pre-training, highlighting the superior ability of our method to preserve pre-trained knowledge.
@@ C.2 Results on SemFPN Head
While Mask2Former [13] is predominantly used as decode head in all DGSS experiments, SoRA is compatible with any decode head. To assess its robustness across different heads, we employ the lightweight SemFPN [44] head to benchmark its performance against Rein [80]. Our experimental results (Tab. 16) indicate that SoRA integrates seamlessly with diverse backbones and heads, consistently surpassing the SOTA baseline.
Appendix D Additional Comparison
@@
In Tables 17, 19, and 19, we present an exhaustive comparison with existing methods to illustrate the broader research landscape across multiple DGSS settings. Additionally, Figures 6, 7, and 8 depict DGSS prediction results on unseen domains for Cityscapes, BDD100k, and Mapillary, respectively, while Fig. 9 provides detection results under various adverse conditions. As evident from the visual comparisons above, SoRA demonstrates remarkable robustness to domain shifts resulting from diverse attributes (e.g., translucency, lighting conditions, road features, geographic variations, weather differences), while also excelling in fine-detail recognition compared to the selected baselines.
@@
Domain generalized recognition requires consistent processing of inputs from diverse domains, whereas domain generalized generation involves generating outputs across a range of domains. Although large-scale Text-to-Image (T2I) models have convincingly demonstrated this ability, it can be compromised in subject personalization tasks involving fine-tuning. As shown in Figures 10 and 11, integrating the SoRA framework in this case enables T2I models to fully leverage their generalization capability to synthesize target subjects in new domains. In summary, our proposed methods effectively facilitate domain-generalizable representation learning by maximally preserving pre-trained knowledge across diverse domains while learning task-specific features.
Appendix E Discussion and Limitations
@@
| SoRA (Ours) | Swin-L | M2F | 56.91 | 51.98 | 60.73 | 56.54 |
@@
| SoRA (Ours) | EVA02-L | M2F | 68.05 | 60.81 | 68.33 | 65.73 |
| SoRA (Ours) | EVA02-L | M2F | 69.94 | 62.48 | 68.33 | 66.92 |
@@
| SoRA (Ours) | DINOv2-L | M2F | 71.82 | 61.31 | 71.67 | 68.27 |
| SoRA (Ours) | DINOv2-L | M2F | 73.63 | 63.33 | 70.98 | 69.31 |
@@
| SoRA (Ours) | DINOv2-L | M2F | 73.16 | 61.90 | 72.73 | 69.26 |
| SoRA (Ours) | DINOv2-L | M2F | 74.85 | 63.59 | 73.92 | 70.79 |
@@
| SoRA (Ours) | DINOv2-L | M2F | 67.02 | 76.45 | 71.74 |
| SoRA (Ours) | DINOv2-L | M2F | 68.08 | 77.87 | 72.98 |

### v2

@@ 5.4 Subject Personalization
Domain generalized recognition requires consistent processing of inputs from diverse domains, whereas domain generalized generation involves generating outputs across a range of domains. Although large-scale Text-to-Image (T2I) models have convincingly demonstrated this ability, it can be compromised in subject personalization tasks involving fine-tuning. DreamBooth (DB) [70] introduces subject personalization, which fine-tunes a pre-trained T2I model on a few reference images to enable it to generate new visual concepts. DB updates all parameters within specific blocks, potentially impairing the pre-trained capability to generate images across various domains. Therefore, we explore whether SoMA can be seamlessly integrated into the DB framework to preserve pre-trained knowledge. For this task, we utilize Stable Diffusion XL (SDXL) [64] for the T2I diffusion model. Fig. 5 presents a visual comparison between DB and SoMA, using identical hyperparameter settings and sample seeds. As seen, SoMA, similar to the original SDXL model, can generate high-quality images of diverse visual domains (e.g., style, texture, pose) and exhibits superior identity preservation compared to DB with prior preservation (p.p.) loss [70]. More results with varied prompts and subjects are provided in the Supplemental.
@@ 6 Conclusion
This paper presents a multifaceted exploration of the distribution of generalizable components in VFMs—examining them at the weight level, block level, and within training dynamics—and introduces the SoMA framework as an effective means to preserve generalizable components. Contrary to recent approaches that passively refine frozen VFM features, SoMA is integrated into all linear layers to effectively facilitate task adaptation, while minimizing interference with generalizable components by initializing the low-rank adapter with minor singular components. Through extensive experiments, we observe that SoMA significantly outperforms the state-of-the-art methods in DGSS, DGOD, and even subject-driven image generation tasks.
Acknowledgement: This work was supported by the National Research Foundation (NRF) grant funded by the Korea government (MSIT) [RS-2025-00562400] and [RS-2022-NR068754].
@@
For a comprehensive understanding of our proposed SoMA framework, we have provided this supplementary material. The following table of contents gives a concise overview and directs readers to specific sections of interest.
@@ A.1 DGSS Settings
The experimental settings for all studies conducted in the main paper are outlined in Tab. 10. Unless otherwise specified, Mask2Former [13] is utilized as the default decode head, and following Rein [80], we adopt only the basic data augmentation used in Mask2Former. Additionally, EMA is selectively employed to ensure stable training. To avoid potential overfitting due to the large model (dimension) size when employing DINOv2-giant as the backbone, we opt to lower the SoMA rank from 16 to 8.
@@ A.2 DGOD Settings
The DGOD settings are detailed in the rightmost two columns of Tab. 10. When applying SoMA to convolution-based backbones such as ResNet [30], we linearize both the patch-level convolution and its weights. Specifically, a single convolution operation can be represented as a linear layer, , where , , and . We then apply SoMA as described in Eq. 1. For ResNet backbones, which possess a much narrower pre-trained knowledge compared to VFMs, we use extensive image corruption techniques to simulate domain shifts, following DivAlign [17]. In contrast, when using DINOv2 [58] as the backbone, we simply utilize basic data augmentation used in Co-DETR [93].
@@ A.3 Subject Personalization Settings
We conduct experiments on the DreamBooth dataset [70], which consists of 30 subjects with 4–6 images per subject. In all experiments the SoMA weights are trained using Adam optimizer for 500 iterations with a learning rate of . We set the adapter rank to and only use a center crop for data augmentation. Inspired by recent findings [25] that the first 10 attention layers of up_blocks.0 in SDXL [64] are pivotal for preserving image content, we fine-tune only these layers. Furthermore, to fully leverage pre-trained image-text joint representations, we freeze the cross-attention modules and apply SoMA solely to the self-attention modules.
@@
Appendix B Detailed Ablations
B.1 Component Analysis
In this subsection, we conduct detailed ablation studies under multiple settings: GTAV Mapillary DGSS and Daytime-Sunny {Dusk-Rainy, Daytime-Foggy} DGOD scenarios. In Tables 12 and 12, we systematically evaluate the effectiveness of each component within the SoMA framework based on class-wise IoU/AP (%). All proposed components enhance overall generalization performance without adding any additional training or inference costs.
As illustrated in Tab. 12, freezing early blocks not only substantially reduces the number of trainable parameters but also significantly improves performance for classes that are infrequently observed in the source dataset (e.g., bicycle, motorcycle, train). Additionally, tuning minor singular components maximizes the retention of VFM’s world knowledge during task adaptation, leading to superior generalization performance over tuning principal components (PiSSA [55]) for most classes. For an in-depth comparison of our methods with PiSSA, please refer to Sec. E. Lastly, annealing weight decay proves especially beneficial for classes requiring fine-detail discrimination (e.g., road vs. sidewalk, traffic light vs. traffic sign, car vs. truck vs. bus vs. train, motorcycle vs. bicycle). Likewise, all components clearly improve recognition for the majority of classes under adverse weather detection settings (see Tab. 12).
@@ B.2 Freezing Scheme
While freezing the initial blocks of VFM is effective in preserving its generalization ability during task adaptation, freezing too many blocks can lead to a reduction in discriminability. To better understand this trade-off, we explore the effects of varying the number of frozen blocks. Tab. 13 shows that freezing up to the first 8 blocks progressively enhances performance, but freezing beyond this point results in a decline. Considering that feature maps from multiple blocks (e.g., the 8th, 12th, 16th, and 24th blocks in large-sized backbones) serve as inputs to the segmentation/detection head, using more than one frozen VFM features as head input significantly undermines task adaptability (i.e. discriminability).
@@ C.1 Results on Various Backbones
Tab. 15 showcases the versatility of SoMA across a wide range of backbones, ranging from isotropic Vision Transformers (ViTs) to ConvNets and hierarchical ViT, as well as models trained under various approaches, such as ImageNet [18] supervision and MAE [31, 81] pre-training. SoMA consistently outperforms FFT across diverse backbone architectures. Notably, the improvements brought by SoMA become increasingly pronounced with larger model sizes and more extensive, high-quality data during pre-training, highlighting the superior ability of our method to preserve pre-trained knowledge.
@@ C.2 Results on SemFPN Head
While Mask2Former [13] is predominantly employed as decode head in all DGSS experiments, SoMA is compatible with any decode head. To assess its robustness across different heads, we employ the lightweight SemFPN [44] head to benchmark its performance against Rein [80]. Our experimental results (Tab. 15) indicate that SoMA integrates seamlessly with diverse backbones and heads, consistently surpassing the SOTA baseline.
Appendix D Model Efficiency Comparison
As detailed in Tab. 17, SoMA exhibits higher throughput than adapter- and VPT-based methods like Rein [80] and SET [86], as it incurs no additional latency. This advantage is especially significant in online inference settings, where the batch size is typically as small as one [33]. Furthermore, in scenarios involving DINOv2-giant exceeding 1B parameters, SoMA can drastically reduce training costs compared to FFT. SoMA initialization is completed within 30 seconds for large-sized models, which is a negligible cost given the improved performance.
Appendix E Additional Comparison
In Tables 18, 20, and 20, we present an exhaustive comparison with existing methods to illustrate the broader research landscape across multiple DGSS settings.
| Efficiency | Training (bs = 4) | Inference (bs = 1 / 32) | ||
| Methods | Time (hrs) | Memory | Throughput (imgs/s) | Memory |
| 9.2 | 12.5G | 20.0 / - | 5.5G / OOM | |
| 9.3 | 12.2G | 33.6 / 64.3 | 4.7G / 48.2G | |
| 9.0 | 12.7G | 56.4 / 79.7 | 4.4G / 40.9G | |
| 27.9 | 45.3G | 21.6 / - | 10.6G / OOM | |
| 18.9 | 25.6G | 21.6 / - | 10.6G / OOM | |
| Synthetic-to-Real Generalization | Test Domains (mIoU in %) | |||||
| Methods | Backbone | Citys. | BDD | Map. | Avg. | |
| Single-source DGSS Trained on GTAV | ||||||
| PEGO [34] | DINOv2-L | 2.6M | 68.86 | 61.44 | 68.61 | 66.30 |
| PiSSAr16 [55] | DINOv2-L | 7.3M | 69.43 | 60.62 | 69.44 | 66.50 |
| SoMA (Ours) | DINOv2-L | 4.9M | 71.82 | 61.31 | 71.67 | 68.27 |
Comparing SoMA with PiSSA [55] and PEGO [34]. PiSSA optimizes parameter efficiency by selectively adjusting the principal singular direction, which is the most stretched direction of the weight matrix. Also PEGO enhances domain generalization by enforcing strict orthogonality between the LoRA adapter and every direction of the pre-trained weights. In stark contrast to PiSSA, SoMA tunes minor singular components, effectively preserving the integrity of pre-trained knowledge. Importantly, our method accounts not only for component orthogonality but also for how pre-trained knowledge is structured within the weight matrix. Thus, although both methods maintain similar orthogonality between tuned and frozen components, SoMA, unlike PiSSA—which directly tunes principal components—robustly adjusts the hierarchical world knowledge structure (see Fig. 2 in the main paper), achieving superior DGSS performance, as shown in Tab. 17. These findings are consistent with the observations presented in the Ablation Sec. B.1.
Furthermore, unlike PEGO, our SoMA leverages spectral information to initialize the LoRA adapter, allowing it to naturally preserve the pre-trained knowledge structure without relying on explicit regularization loss. As evidenced by its superior performance relative to PEGO in Tab. 17, we argue that imposing strict orthogonality constraints on all directions of the pre-trained weights may excessively restrict task adaptation, potentially compromising discriminability. Lastly, whereas PEGO and PiSSA explore adaptation solely at the weight level, our SoMA framework extends its analysis to both the block level and training dynamics.
@@ Appendix E Additional Comparison
DGSS and DGOD Qualitative Comparison. Figures 6, 7, and 8 depict DGSS prediction results on unseen domains for Cityscapes, BDD100k, and Mapillary, respectively, while Fig. 9 provides detection results under various adverse conditions. As evident from the visual comparisons above, SoMA demonstrates remarkable robustness to domain shifts resulting from diverse attributes (e.g., translucency, lighting conditions, road features, geographic variations, weather differences), while also excelling in fine-detail recognition compared to the selected baselines.
@@
Subject Personalization. Domain generalized recognition requires consistent processing of inputs from diverse domains, whereas domain generalized generation involves generating outputs across a range of domains. Although large-scale Text-to-Image (T2I) models have convincingly demonstrated this ability, it can be compromised in subject personalization tasks involving fine-tuning. As shown in Figures 10 and 11, integrating the SoMA framework in this case enables T2I models to fully leverage their generalization capability to synthesize target subjects in new domains.
In summary, our proposed methods effectively facilitate domain-generalizable representation learning by maximally preserving pre-trained knowledge across diverse domains while learning task-specific features.
Appendix F Discussion and Limitations
@@
| SoMA (Ours) | Swin-L | M2F | 56.91 | 51.98 | 60.73 | 56.54 |
@@
| SoMA (Ours) | EVA02-L | M2F | 68.05 | 60.81 | 68.33 | 65.73 |
| SoMA (Ours) | EVA02-L | M2F | 69.94 | 62.48 | 68.33 | 66.92 |
@@
| SoMA (Ours) | DINOv2-L | M2F | 71.82 | 61.31 | 71.67 | 68.27 |
| SoMA (Ours) | DINOv2-L | M2F | 73.63 | 63.33 | 70.98 | 69.31 |
@@
| SoMA (Ours) | DINOv2-L | M2F | 73.16 | 61.90 | 72.73 | 69.26 |
| SoMA (Ours) | DINOv2-L | M2F | 74.85 | 63.59 | 73.92 | 70.79 |
@@
| SoMA (Ours) | DINOv2-L | M2F | 67.02 | 76.45 | 71.74 |
| SoMA (Ours) | DINOv2-L | M2F | 68.08 | 77.87 | 72.98 |